### v1

# Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1

Abstract
[…]
*[Uncaptioned image]DDroid: A Resource Suite for AI-Generated Code Detection
Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2 Ubiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technische Universität Darmstadt {name.surname}@mbzuai.ac.ae; {name.surname}@tu-darmstadt.de
1 Introduction
In recent years, language models (LMs) for code generation (Code-LMs) have become a near-indispensable accessory in a developer’s toolbox. Their enhancement of productivity has proliferated into most of the software development lifecycle, including automating unit test generation (Jain et al., 2025), code infilling (Bavarian et al., 2022), predicting build errors, and code refactoring, inter alia, propelling their broad adoption in production (Dunay et al., 2024; Frömmgen et al., 2024; Murali et al., 2024). However, the code authoring and refinement abilities of these models present issues with respect to domains where the human authorship of the generated artefacts is paramount and the consequences of limited human supervision are of concern.
Despite the well-documented productivity benefits of using AI assistance for knowledge workers (Weber et al., 2024b; Li et al., 2023a), there exists a wide range of scenarios where ensuring the human authorship of artefacts is vital, resulting in the need for robust detectors of machine-assisted code. For instance, in academia, students’ reliance on LMs for assignments undermines educational integrity, with professors unable to detect the authorship of submissions and grading AI-generated coding practices (Koike et al., 2024). Similarly, conducting technical hiring fairly and human code annotation studies accurately requires the ability to ensure that the submitted artefacts are authentically human-authored (Veselovsky et al., 2023).
The subtle failure patterns in the outputs of code LMs imply the need for strong detection mechanisms as part of the workflow in order to safeguard against unforeseen side effects. For instance, machine-generated code can introduce serious vulnerabilities (e.g., insecure logic, hidden backdoors, or injection flaws), which can jeopardise software reliability (Bukhari, 2024) and data security (Pearce et al., 2025). It can also facilitate obfuscation, producing code that is harder to parse (Vaithilingam et al., 2022): this can hide malicious functionality and complicate debugging (Nunes et al., 2025). These weaknesses can amplify over time, creating a dangerous feedback loop where (possibly deficient) AI-generated code enters public repositories and is leveraged for subsequent training runs, thus increasing the risk of degraded data quality (Ji et al., 2024) or, even worse, collapsing models (Shumailov et al., 2024).
Despite the increasing interest in detecting AI-generated code, most current work has notable limitations. Existing work usually covers fewer than three programming languages (Xu et al., 2025) and focuses on a narrow set of API-based code generators (Yang et al., 2023). Moreover, detectors typically address the problem as a binary classification task: machine-generated vs. human-written code (Jawahar et al., 2020). This ignores common hybrid operating modes where code is co-authored by humans and LMs or adversarial scenarios where models are prompted or tuned to evade detection (Abassy et al., 2024).
Our work addresses these limitations with a comprehensive and scalable approach to AI-generated code detection. Our contributions are as follows:
  • We compile and open-source DroidCollection, an extensive suite of multi-way classification data for training and evaluating AI-generated code detectors. DroidCollection contains over 1 million instances sampled from 43 LMs (spanning 11 model families), 7 programming languages, and multiple coding domains.
  • We propose a novel task: detection of code generated by adversarially trained LMs, which mimics intentional obfuscation and evasion behaviours. To this end, we compile and release DroidCollection-Pref, a preference corpus of 157k response pairs designed to encourage language models to produce responses that closely resemble those of humans.
  • We open-source DroidDetect-Base and DroidDetect-Large, two state-of-the-art AI-generated code detectors fine-tuned from ModernBERT (Warner et al., 2024) Base (149M), and Large (396M) models, respectively, using DroidCollection.
  • We conduct extensive out-of-distribution performance analysis across languages, coding domains and detection settings. Our evaluation results demonstrate that there is positive transfer across related programming languages (Martini, 2015) and across domains. We also find that most existing models struggle when tasked with detecting machine-refined code and are almost entirely unusable against adversarially humanised model-generated code. However, we show that this can be rectified by incorporating modest amounts of such data during training.
2 Related Work
We briefly outline three relevant lines of existing work: 1) AI-generated text detection, 2) AI-generated code detection, and 3) adversarial evasion of AI-generated content detectors.
2.1 AI-Generated Text Detection
Early research on synthetic data detection has focused on detecting AI-generated text in specific, fundamental tasks such as question answering (Guo et al., 2023), translation, summarisation and paraphrasing (Su et al., 2023). Major early contributions to creating comprehensive benchmarks include M4 (Wang et al., 2024), which introduced a multilingual, multi-generator and multi-domain benchmark consisting of 122,000 human-written and machine-generated texts. MULTITuDE (Macko et al., 2023) featured a multilingual dataset with over 70,000 samples of AI and human-written texts across 11 languages. Additionally, MAGE (Li et al., 2024) concentrated on English-only scenarios, but emphasised evaluating model robustness by testing across eight distinct out-of-domain settings to simulate real-world scenarios. The advancement of this field has been further stimulated by numerous competitions and shared tasks dedicated to AI-generated text detection, including RuATD (Shamardina et al., 2022), a shared task at COLING’2025 (Wang et al., 2025), a shared task at ALTA (Mollá et al., 2024), and DagPap (Chamezopoulos et al., 2024).
Tools such as MOSS (Puryear and Sprint, 2022) have shown some effectiveness in identifying AI-generated code, since their style is out of the ordinary distributions of student solutions. However, Pan et al. (2024) and JianWang et al. (2024) have shown that detectors such as GPT-Zero often fail when applied to code rather than text. This critical observation, backed up by our experiments in Section 4, highlights the inadequacy of directly porting generic text-based models to the code domain and strongly motivates the creation of code-specific detection strategies and specialised datasets. Our work responds to this need by providing a large-scale, multifaceted suite specifically curated for AI-generated code, designed to foster the development and rigorous testing of detection techniques attuned to the unique characteristics of programming languages and AI-generated software.
2.2 AI-Generated Code Detection
Early attempts at AI-generated code detection using decision tree learning methods, such as Idialu et al. (2024) and Li et al. (2023b), demonstrated that code-level statistics (e.g., number of lines, Abstract Syntax Tree (AST) depth, identifier length) can serve as reliable indicators of authorship. However, more sophisticated detection requires a more involved feature engineering, which is best performed using deep learning methods (Tulchinskii et al., 2023). Thus, recent efforts have mainly focused on training text-based LMs to detect AI-generated code. A common approach in existing work, such as GPTSniffer (Nguyen et al., 2024) and GPT-Sensor (Xu et al., 2025) is to extract human-written functions from the CodeSearchNet dataset (Husain et al., 2020) and then to prepare machine-generated counterparts to them using ChatGPT. Although similar in their dataset construction, these two works differ in modelling. GPTSniffer uses a multi-class classification loss, whereas GPT-Sensor applies a cosine similarity-based loss to better separate the embeddings of AI-generated and human-written code, aiming at learning more discriminative representations.
To address the overdependence on CodeSearchNet in prior work, Orel et al. (2025) source additional code from LeetCode and CodeForces. They evaluated a wide range of locally deployable LMs as code generators and provided a systematic analysis of out-of-distribution (OOD) detection performance across different settings. Importantly, they go beyond binary classification by introducing more nuanced scenarios, such as collaborative (or hybrid) settings where LMs complete or rewrite human-written programs.
CodeMirage Guo et al. (2025), another benchmark released concurrently with ours, also includes rewritten code. However, these works lack diversity in terms of generators: CoDet-M4 covers only 5 generators, while CodeMirage covers 10. Furthermore, they overlook the importance of diverse sampling strategies in the detection of machine-generated codes and do not consider more adversarial settings where the model-generated code is artificially humanised. Our work builds upon and extends the progress of previous works by further increasing the scale and diversity: we incorporate three distinct domains – competitive programming solutions (e.g., LeetCode), open-source GitHub repositories, and code from research papers – utilise 43 generative models, and cover seven programming languages. Notably, unlike Codet-M4 or CodeMirage, our dataset is the first in this domain to systematically integrate diverse sampling strategies using varied generation settings and synthetic data generation scenarios.
2.3 Adversarial Evasion of AI-Generated Content Detectors
Although specialized detectors for AI-generated code can be effective against honest actors, their straightforward training on machine-generated and machine-refined data makes them vulnerable to adversarially perturbed or humanised text, modified to evade detection (He et al., 2024; Masrour et al., 2025). Currently, RAID (Dugan et al., 2024), one of the most extensive benchmarks in AI-generated text detection, is notable in being one of the few efforts exploring adversarial detection settings with various attack methods such as paraphrasing and synonym substitution. Our work in AI-generated code detection builds upon this important aspect. We extend this focus to the code modality by incorporating a diverse set of adversarial attack scenarios specifically engineered to challenge detectors. Moreover, we move beyond the language manipulations considered by RAID to address the possibilities of adversarial training using targeted mining of paired preference data and a dedicated collection of adversarial prompting, which are all aspects that are vital for assessing detector robustness under more challenging conditions.
3 The DroidCollection Corpus
Name Size Supported No. of Supported Varied Machine Adversarially
Domains Models Languages Sampling Refined Data Humanized Data
GPT-Sniffer (Nguyen et al., 2024) 7.4k 1 1 2
CodeGPTSensor (Xu et al., 2025) 1.1M 1 1 2
Whodunit (Idialu et al., 2024) 1.6k 1 1 1
CoDet-M4 (Orel et al., 2025) 501K 2 5 3
CodeMirage (Guo et al., 2025) 210k 1 10 10
DroidCollection 1.06M 3 43 7
Table 1: Comparison of DroidCollection to other AI-generated code detection datasets shows broader domain and model coverage, as well as unique inclusion of varied sampling strategies and adversarially humanized data.
We compare our dataset with existing ones in Table 1, which shows that our dataset is not only among the largest to date, but also captures a broader range of variations. Additional details about key characteristics are provided in Section B.3.
In this section, we detail the curation of the human-generated, machine-generated, and machine-refined splits of DroidCollection. The adversarially humanised data collection is deferred to Section 3.4.
3.1 Human-Authored Code Acquisition
To build the dataset, we collected human-written samples from multiple sources, covering C++/C, C#, Go, Java, JavaScript and Python languages. Then, we generated code, using base and instruction-tuned LMs from 11 model release families, namely: Llama (Grattafiori et al., 2024), CodeLlama, GPT-4o, Qwen (Qwen et al., 2025), IBM Granite (Mishra et al., 2024), Yi (AI et al., 2025), DeepSeek (Guo et al., 2024), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), Mistral (AI, 2025), Starcoder (Li et al., 2023c). The list of generators per model family is detailed in Appendix A. Our dataset covers three domains: general use code, algorithmic problems, and research/data-science code.
General Use Code
This type of code is normally deployed for disparate use cases such as web serving, firmware, game engines, etc. These are largely hosted on GitHub, and mainly obtained from StarcoderData (Li et al., 2023c), and The Vault (Manh et al., 2023) datasets.
Algorithmic Problems
This category contains code solutions to competitive programming problems – algorithmic challenges designed to test problem-solving skills, commonly featured in competitions. These are retrieved from multiple sources such as TACO (Li et al., 2023d), CodeNet (Puri et al., 2021) (mainly AtCoder [^2] https://atcoder.jp/ and AIZU [^3] https://onlinejudge.u-aizu.ac.jp/home platforms), LeetCode and CodeForces, from the work of  Orel et al. (2025). Its primary distinguishing feature is its tendency to contain simple and self-contained routines.
Research Code
This data subset is sourced from code repositories accompanying research papers, collected in ObscuraCoder (Paul et al., 2025). Additionally, we augment this data with mathematical and data science code from work of  Lu et al. (2025). This subset is characterised by the lack of modularity and the over-representation of procedural code.
3.2 AI-Authored Code Generation
Generation via Inverse Instruction
Since the data from sources such as CodeNet and StarcoderData do not contain any instructions, we decided to apply inverse instructions to transform code from these datasets into instructions, which can be used to prompt LMs. In our case, the method of preparing inverse instructions was similar to that described in InverseCoder (Wu et al., 2025): we passed the code snippets to an LM, asking it to build a summary, and a step-by-step instruction that can be given to an LM to generate a similar code. The main difference between our approach and that of InverseCoder is that we sought to minimise the costs of the generation and did not split the summarisation and instruction generation into separate LM calls. However, in cases where a summary could be extracted from the response but the instruction could not, we used the summary to re-generate the instruction. This experiment with details about the prompts and the models we used is illustrated in Section B.1. This type of generation allows us to cover a wide range of prompts, simulating a diversity of user-LM interactions, which is common in the real world.
Generation Based on Comments
Some of the data sources used in our study provide docstrings (The Vault Class and Function) or comments (The Vault Inline) that describe the given code. In this case, we mainly used base models, which were prompted with the first line of code and the docstring or comment for generation. Instruct models were given only the docstring and a task to implement the desired class or function.
Generation Based on a Task
The examples from platforms with algorithmic problems mainly come with a precise task description or a problem statement. In this case, we only used the description to prompt the LMs for generation.
Unconditional Synthetic Data
This data is not conditioned on prior human generations. The rationale behind this is that the machine-generated data used to train the majority of AI-generated content detectors is acquired in a biased manner. It usually involves completing incomplete human-written samples or responding to prompts conditioned on existing human generations. This bias, though rather subtle, leads to a situation where detectors are only exposed to the kinds of synthetic data that are easiest for the models to learn (Su et al., 2024). Hence, we seek to obtain synthetic data that is not conditioned on prior human generations. Following prior work [^4] https://huggingface.co/blog/cosmopedia, we create synthetic user profiles on which we condition coding tasks and, in turn, the final generated code.
To explore how large language models can simulate the behaviour profiles of real programmers, we took inspiration from the PersonaHub dataset (Ge et al., 2024). We first generated a diverse set of programmer profiles, and then used an LM to create programming tasks that can typically be performed by programmers of such types. These tasks, along with their corresponding descriptions, termed DroidCollection-Personas, were then used to generate code samples. More details about the DroidCollection-Personas are outlined in the Section B.2.
3.3 Machine-Refined Data
In practice, purely AI-generated code is rare. Developers typically collaborate with LMs, starting with human-written code and asking the model to modify or extend it. This makes binary classification (human vs. machine) insufficient for real-world scenarios. Instead, introducing a third class to capture human-LLM collaboration, as proposed by Orel et al. (2025), offers a more realistic and useful approach.
To generate such samples, we designed three scenarios: (i) Human-to-LLM continuation: A human initiates the code, and the LM completes it. We simulated this by preserving the first $N\%$ of the code lines (where $N$ ranges from 10% to 85%) and asking the model to complete the rest. (ii) Gap filling: The model is given the beginning and the end of a human-written code snippet and is asked to generate the missing part in the middle. The amount of preserved code follows a similar proportion as in the first scenario. (iii) Code rewriting: The LM is asked to rewrite human-authored code, either with no specific prompt or with an instruction to optimise it.
3.4 Adversarial Samples
With the development of post-training techniques such as PPO (Schulman et al., 2017), DPO (Rafailov et al., 2023), and GRPO  (Shao et al., 2024), it has become possible to set up the training in adversarial ways that enable LMs to evade AI-generated code detectors. Prior work by Shi et al. (2024) and Sadasivan et al. (2023) has shown that LM-generated content detectors are vulnerable to adversarial attacks and spoofing. This motivates the inclusion of adversarial samples in DroidCollection to improve model robustness.
To this end, we introduce two types of adversarial attacks: prompt-based attacks and preference-tuning-based attacks. In the prompt-based setting, we construct adversarial prompts by instructing the model to “write like a human” in multiple ways, relying on the models’ parametric knowledge of how to produce outputs that mimic human-authored code and thus challenge detection systems. In the preference-tuning-based setting, we curate DroidCollection-Pref, a dataset of 157K paired examples consisting of human-written and LM-generated code responses to the same prompt. Using DroidCollection-Pref, we train LMs with up to 9B parameters –including LLaMA, Qwen, and Yi–, using LoRA (Hu et al., 2022) with rank 128 and DPO for two epochs. These models’ output distributions are, in effect, steered towards preferring human-like code, making them less likely to contain the stylistic features of machine-generated code. Once trained, the models are used to generate new “machine-humanised” code samples. We filter their outputs as in Section 3.6 to keep only high-quality adversarial examples. Finally, we obtained a nearly 1:1 ratio of prompt-based vs. preference-tuning adversarial attacks.
3.5 Varying Decoding Strategies
Strategy Attribute Range
Greedy Decoding
Sampling Temperature {0.1, 0.4, 0.7, 1.0, 1.5, 2.0}
Top-k {10, 50, 100}
Top-p {1.0, 0.95, 0.9, 0.8}
Beam Search Beam Width {2, 4, 6, 8}
Table 2: Decoding settings used for the AI-generated, AI-refined, and AI-humanised splits of DroidCollection.
It was shown by Ippolito et al. (2020) that after greedy decoding, it was easier to detect AI-generated texts compared to when using other decoding techniques. To capture this challenging behaviour and reflect the diversity of generative systems, we experimented with various decoding strategies, as shown in Table 2.
3.6 Data Filtering
To ensure the quality of our DroidCollection dataset, we applied a series of filtering criteria, commonly used in other code-related works (Lozhkov et al., ; Li et al., 2023c; Paul et al., 2025). First, we removed code samples that could not be successfully parsed into an abstract-syntax tree (AST). We also filtered samples based on the AST depth, keeping only those with the depth between 2 and 31, to avoid too simple or too complex codes. We restricted each sample’s maximum line length to be between 12 and 400 characters, and the average line length to fall between 5 and 140 characters, and used only samples with between 6 and 300 lines of code. Moreover, we filtered samples according to the fraction of alphanumeric characters, retaining only those between 0.2 and 0.75, to avoid the usage of configs and auto-generated files. To ensure English documentation, we used the Lingua language detector [^5] GitHub: pemistahl/lingua-py and retained only samples where the docstrings showed greater than 99% confidence of being English. Finally, we removed duplicate or near-duplicate samples; for this, we used MinHash (Broder, 1997), with a shingle size of 8 and a similarity threshold of 0.8.
It should be acknowledged that we did not filter out the human-written codes which were sourced after the coding copilots became popular. It means, that there is a possibility that among human written codes there could be samples, created with the help of LLMs. This potential issue is tackled in Section 5.
Model 2-Class 3-Class
General Algorithmic Research/DS Avg. General Algorithmic Research/DS Avg.
Zero-Shot Baselines Fast-DetectGPT (Bao et al., 2024) 75.07 63.05 65.43 67.85 66.43 62.90 64.30 64.54
CoDet-M4 (Orel et al., 2025) 53.41 44.63 65.43 54.49 41.90 46.06 55.43 47.80
M4 (Wang et al., 2024) 50.17 57.91 44.67 50.92 56.46 58.13 51.21 55.27
GPTSniffer (Nguyen et al., 2024) 54.25 36.85 32.10 41.07 45.22 31.75 39.88 38.95
GPTZero 54.05 71.96 44.73 56.91 50.56 66.13 30.62 49.10
OOD Evaluation DroidDetectCLS-BaseGeneral 99.30 53.73 76.46 76.50 93.05 46.22 76.99 72.09
DroidDetectCLS-BaseAlgorithmic 49.63 98.26 60.78 69.56 47.86 92.84 56.58 65.76
DroidDetectCLS-BaseResearch/DS 47.01 48.02 72.55 55.86 47.86 38.73 59.97 48.85
Fine-Tuned Baselines GCN 78.57 60.61 67.79 68.99 56.85 46.91 51.13 51.63
CatBoost 89.69 87.29 77.21 84.73 78.86 74.01 64.07 72.31
Full Training M4FT 92.99 89.36 73.99 85.45 80.98 80.72 58.80 73.50
GPT-SnifferFT 97.72 96.52 80.46 91.56 89.42 88.12 70.72 82.75
CoDet-M4FT 98.89 98.23 83.77 93.63 85.46 90.41 73.88 83.25
DroidDetectCLS-Base 99.22 98.22 87.57 95.00 92.78 93.05 74.46 86.76
DroidDetectCLS-Large 99.38 98.39 93.24 97.00 93.08 92.86 80.42 88.78
Table 3: Comparison of models in 2-Class (human- vs machine-generated) and 3-Class (human- vs machine-generated vs machine-refined) classification setups across programming languages in terms of weighted F1-score. In the OOD section, we show models trained on each domain individually. FT subscript in Full Training section means that this model was fine-tuned on DroidCollection The best results are shown in bold.
4 Detection Experiments
4.1 Experimental Setup
We first evaluate a range of detectors to assess the strengths and limitations of existing AI-generated code detection models. We include several off-the-shelf detectors as our zero-shot baselines and models trained on DroidCollection as fine-tuned baselines.
The baselines include models widely used in related works: (i) GPT-Sniffer(Nguyen et al., 2024), a CodeBERT-based binary classifier fo AI-Generated code detection; (ii) CoDet-M4(Orel et al., 2025), a UnixCoder model trained on outputs from multiple code generators; (iii) M4 classifier(Wang et al., 2024), for general AI-generated text detection; (iv) Fast-DetectGPT(Bao et al., 2024), a distribution-based zero-shot detector; and (v) GPT-Zero [^6] https://gptzero.me/, an API-based detector (as this API is paid, we evaluated it on a representative sample of 500 code snippets for each label-language and label-domain pair). Since most of these baselines use binary classification, we map our ternary labels (human-written, AI-generated, AI-refined) into binary targets for fair comparison.
Additionally, we train other models directly on our dataset using a multi-class objective: (i) a simple GCN(Kipf and Welling, 2017); (ii) a CatBoost classifier(Prokhorenkova et al., 2018), following procedures similar to the Whodunit paper; and (iii) two encoder-only transformers, ModernBERT-Base and ModernBERT-Large (Warner et al., 2024), denoted as DroidDetectCLS-Base and DroidDetectCLS-Large. Full details are provided in Sections C.1, C.2 and 5.
4.2 RQ1: What is the value of extensive data collection for training robust detectors?
Table 1 highlights a key limitation of existing datasets: they often lack diversity. Tables 3 and 4 show that this limitation significantly impacts detector performance in realistic settings. Baseline detectors - the ones illustrated in Zero-Shot Baselines section of the tables - underperform on our test split compared to even simple fine-tuned baselines like GCN and CatBoost. Among the baselines evaluated, zero-shot Fast-DetectGPT consistently yields strong performance across both languages and domains, outperforming all other baselines. In contrast, pre-trained models usually perform well only on languages and domains that are closely aligned with their original training data. This highlights the limitations of previously collected datasets, which do not cover the diversity of generations in DroidCollection, and hence are far from being useful in real-life scenarios.
The Full Training section of tables contains the variations of pre-trained models from Zero-Shot Baselines, but fine-tuned on DroidCollection. It is clearly seen, that after training on DroidCollection data, the performance of the models significantly increases. They start outperforming GCN and CatBoost, but still lag behind DroidDetectCLS models.
DroidDetectCLS models trained on our extensive training split consistently outperform all baselines, achieving near-ideal scores in both binary and ternary tasks. Notably, the larger backbone (DroidDetectCLS-Large) dominates across all settings, demonstrating that both model size and diverse training data are crucial for high classification performance.
4.3 RQ2: How well do models generalise in OOD settings?
We evaluated the DroidDetectCLS-Base backbone in OOD settings under language shift and domain shift conditions: training it on a single programming language or domain. Comparing the multi-domain and the multi-lingual performance of the baselines to our backbone models trained in such restricted conditions allows us (i) to uncover possible shortcomings in the training data curation of popular baseline models, as they can be compared head-to-head to both the split-specific and fully-trained variants of our backbone, and (ii) to assess the inherent ease with which models are able to transfer along these settings, by-proxy outlining the value of extensive training data curation. We selected the base version of the backbone for this restricted training scenario since it was comparable to most of the chosen baselines in terms of size.
Table 4 shows that, under restricted training conditions, models tend to generalise better to syntactically similar languages. For example, a model trained on C/C++ performs reasonably well on C# and Java. However, for typologically isolated languages such as Python or JavaScript, all models not trained specifically for it tend to struggle in this setting. Table 3 illustrates that the models trained on a single domain have a high discrepancy in scores for other domains. For example, as can be seen in the OOD Evaluation section, the classifiers trained on algorithmic problems suffer with the general codes, and show comparable low performance on Research/DS codes, and vice versa.
4.4 RQ3: How robust are models to adversarial samples?
Finally, we test adversarial robustness using challenging samples designed to evade detection (Table 5). Here, many baseline detectors struggle: for instance, GPT-Zero achieves only 0.10 recall. M4 and CoDet-M4 show higher recall on adversarial samples but also have high false positives, misclassifying human-written code. Interestingly, fine-tuning on DroidCollection also improves the recall of baselines for human-written cases. Remarkably, the recall of M4 on adversarial cases drops from 0.73 to 0.67, while the recall on human-written cases increases significantly from 0.40 to 0.91. A similar pattern is observed for CoDet-M4, while for GPT-Sniffer, both scores increase.
In contrast, DroidDetectCLS-Base, trained with explicit exposure to such samples, maintains strong performance with a recall above 0.9. This shows that training on diverse, adversarially crafted examples further enhances robustness and reduces susceptibility to trivial evasion strategies.
Model 2-Class 3-Class
C/C++ C# Go Java Python JS Avg. C/C++ C# Go Java Python JS Avg.
Zero-Shot Baselines Fast-DetectGPT (Bao et al., 2024) 81.33 72.77 81.16 76.03 73.60 74.59 76.58 77.85 66.37 72.73 69.45 70.34 69.11 70.98
CoDet-M4 (Orel et al., 2025) 61.12 50.68 19.66 56.15 58.75 41.44 47.97 53.81 40.74 18.28 45.26 53.51 36.09 41.28
M4 (Wang et al., 2024) 62.22 40.73 57.59 48.39 61.47 64.44 52.81 65.33 50.38 60.49 56.25 61.21 53.64 57.92
GPTSniffer (Nguyen et al., 2024) 63.02 48.90 79.89 40.30 38.34 45.94 52.40 64.18 42.29 76.19 34.94 34.94 47.22 49.96
GPTZero 58.32 45.69 13.64 74.65 73.19 63.16 54.81 61.00 50.38 28.89 61.25 52.63 54.78 51.48
OOD Evaluations DroidDetectCLS-BaseC/C++ 98.98 96.59 67.32 96.97 74.45 91.15 87.58 92.62 81.67 56.43 79.45 56.43 69.72 72.72
DroidDetectCLS-BaseC# 93.66 99.20 78.89 95.20 71.13 89.87 87.99 80.95 92.93 57.74 84.17 54.25 65.18 71.04
DroidDetectCLS-BaseGo 93.33 86.00 98.94 89.97 71.45 88.72 88.07 80.74 63.61 92.93 74.18 50.38 65.37 71.20
DroidDetectCLS-BaseJava 95.53 96.42 94.57 99.31 75.59 80.26 90.28 85.00 84.43 58.85 93.38 63.25 64.57 74.91
DroidDetectCLS-BasePython 80.27 85.48 82.28 88.80 98.85 86.62 86.75 67.59 75.56 53.70 79.31 93.08 69.96 73.20
DroidDetectCLS-BaseJS 95.76 97.38 75.27 96.45 68.98 97.80 88.61 87.96 87.58 52.78 86.32 60.78 89.67 77.52
Fine-Tuned Baselines GCN 79.06 78.33 84.33 80.04 72.49 69.69 77.32 65.97 58.03 65.20 60.13 55.22 54.72 59.88
CatBoost 94.00 91.20 90.57 92.26 89.51 82.55 90.02 84.57 81.32 81.54 82.42 78.15 70.98 78.83
Full Training M4FT 94.18 89.98 92.19 92.63 87.19 93.61 91.63 79.56 75.55 77.63 79.55 69.63 77.53 76.57
GPT-SnifferFT 97.64 97.36 97.33 97.96 95.07 97.94 97.22 85.14 85.75 85.78 86.97 79.02 88.28 85.16
CoDet-M4FT 99.36 99.22 99.31 99.04 98.28 99.24 99.08 89.98 88.94 89.73 91.70 85.80 91.46 89.60
DroidDetectCLS-Base 99.29 99.33 99.32 99.45 98.87 98.38 99.11 94.43 94.06 93.98 93.93 93.95 90.99 93.56
DroidDetectCLS-Large 99.31 99.51 99.32 99.45 99.11 98.67 99.23 94.24 93.87 94.42 94.05 94.13 91.27 93.66
Table 4: Comparison of models in 2-class (human- vs. machine-generated) vs. 3-class (human- vs. machine-generated vs. machine-refined) classification setups across programming languages in terms of weighted F1-score. In the OOD section, we train on each programming language individually. The best results are highlighted in bold.
FastDetectGPT GPTSniffer M4 CoDet-M4 GPT-Zero M4FT GPT-SnifferFT CoDet-M4FT DroidDetectCLS-Base DroidDetectCLS-Large
Human-written 0.84 0.65 0.40 0.38 0.53 0.91 0.97 0.96 0.93 0.98
Adversarial samples 0.48 0.49 0.73 0.63 0.10 0.67 0.55 0.51 0.92 0.92
Table 5: Recall for human-written vs. adversarial examples. The red cells show that despite having high recall on adversarial samples, M4 and CoDet-M4 struggle to detect human-written code. The best results are in bold.
5 Detector Training and Ablations
We conducted a series of ablation experiments starting from our DroidDetectCLS backbone to systematically identify the most effective model architecture and training strategy.
As an architectural ablation, we explore whether incorporating the structural representation of code could improve the detector’s performance. Specifically, we trained a 4-layer Graph Convolutional Network over the AST representation of codes to evaluate its ability to distinguish AI-generated from human-written code. The results are shown in Section C.1. We can see that while structural signals are informative, GCNs alone are not sufficient to achieve strong generalisation.
Next, we explored early fusion of textual and structural representations by combining a text encoder with a GCN encoder. For text encoding, we used ModernBERT (Warner et al., 2024), a transformer-based model pre-trained on a mixture of natural language and code. We experimented with both the base (149M parameters) and the large (396M parameters) variants. This model was selected for the inference efficiency (Warner et al., 2024), and suitability for code-related tasks. However, as shown in Section C.3, this fusion strategy yielded only a marginal improvement. Consequently, we decided to use a text-only encoder for the final model.
We then address the issue of class separability, which can arise because adversarial and refined code is similar to human-written code. We explore training our models using triplet loss (Hoffer and Ailon, 2015) in a supervised contrastive (Khosla et al., 2020) setup using the class labels. This metric-learning approach encourages the model to place samples of the same class closer to each other in the embedding space while pushing dissimilar samples apart, and it has been demonstrably effective in other detection scenarios that require high precision (Deng et al., 2022; Li and Li, 2024). We refer to these models as DroidDetectSCL. Table 6 demonstrates the small but consistent performance gain unlocked using metric learning.
Finally, we addressed the problem of noisy and mislabelled training data. Despite extensive data filtering, it is possible that some code samples curated as human-written may have been generated by coding-copilots. The presence of such examples could negatively impact the training of our detector. To address this, we applied MC Dropout (Hasan et al., 2022) to estimate the model uncertainty on the human-written portion of the test set. Specifically, we identified the top 7% most uncertain samples– those for which a pre-trained model exhibited low prediction confidence–and resampled the dataset, removing them from the training set. We then retrained the model on the remaining data, thereby getting rid of the influence of potentially mislabelled or ambiguous samples. This manner of self-bootstrapping datasets has an extensive track record in image (Yalniz et al., 2019; Xie et al., 2020) and text (Wang et al., 2022) representation learning, relying on the tendency of neural networks to understand patterns in clean labels before they overfit to noisy data (Feldman and Zhang, 2020). Incorporating this filtering into our training yielded our final DroidDetect models, which, as shown in Table 6, perform the best across the classification settings.
We trained all models for 3 epochs, using AdamW (Loshchilov and Hutter, 2019) optimiser, setting the top learning rate to 5e-5, and applying the linear warmup (proportion 0.1) with cosine decay learning rate scheduler. The batch size is 64 for DroidDetect-Base and 40 for DroidDetect-Large.
Model Variant 2-class 3-class 4-class
Base Large Base Large Base Large
DroidDetect 99.18 99.25 94.36 95.17 92.95 94.30
- Resampling 99.15 99.22 93.86 94.43 92.52 93.14
      [DroidDetectSCL]
- Triplet Loss 99.14 99.18 90.51 94.07 89.63 92.65
      [DroidDetectCLS]
Table 6: Weighted F1-score for DroidDetect across training ablations. The best results are shown in bold.
6 Conclusion and Future Work
We have curated DroidCollection, a large and diverse suite of datasets that facilitate the training and the evaluation of robust AI-generated code detectors to support their most common modes of operation, i.e., completion and rewriting, along with potentially adversarial use cases. Among openly available corpora for training AI-generated content detectors, DroidCollection offers the most exhaustive coverage with respect to the number of generators, generation settings, programming languages, and domains covered. We further developed DroidDetect, a suite of AI-assisted code detection models, in two sizes (Base and Large), and conducted extensive ablation studies to evaluate which training strategies yield the most effective performance for this task.
In future work, we plan to enhance the coverage of DroidCollection and the robustness of DroidDetect. We will incorporate code samples from more closed-source API-based generators, broadening the diversity of the code samples. We will also add generations from reasoning LMs to enhance the applicability of our detectors. Finally, we plan to expand language coverage to include languages such as PHP, Rust, and Ruby, thereby making our benchmarks more representative.
Limitations
Corpus Updates and Coverage
Possessing a perfect coverage over all major models in the current fast-paced release environment is an intractable task. We acknowledge that the release of new model families with unseen output distributions presents a challenge for all AI-generated content detectors. Since we have mature pipelines for machine-generated, machine-refined and adversarially-humanised data acquisition, we plan to update DroidCollection with generations sourced from future model releases.
Cost Effectiveness
Owing to cost realities, the majority of training samples in our study are sourced from locally deployable models. The high costs of API invocations are the primary reason why our study leaves data collection from recently released reasoning/thinking models such as Anthropic’s Claude 3.7, DeepSeek R1, and Google’s Gemini 2.5 for future work. For similar reasons, our evaluation of API-based detectors such as GPTZero was limited to a subset of the test set.
Potential Data Contamination
In spite of the thorough curation and extensive filtering undertaken for DroidCollection, we acknowledge the possibility that a small number of AI-generated or AI-assisted code samples may still be mislabeled as human-authored, due to the inherent nature of the data sources used for the dataset construction. Seeking to limit the negative effects of mislabeled or noisy data, our work explores uncertainty-based dataset re-sampling using a pre-trained classifier, which we show to be effective in improving the model’s performance by identifying ambiguous samples to discard during training. In the released dataset, we include flags for code snippets identified as suspicious, enabling downstream users to apply additional filtering or analysis as needed.
Ethics Statement
The human-written code samples in our dataset are sourced exclusively from publicly available code corpora vetted for appropriate licensing and PII removal. Additionally, all code generation was conducted in compliance with the terms of use of the respective model providers.
DroidDetect and DroidCollection aim to promote transparency in code authorship, especially in academic and research settings. While there is a risk that they could be misused to train models to evade detection, we strongly discourage any malicious or privacy-invasive applications. We advocate for the responsible use in strictly legitimate research and educational contexts.
Appendix A List Of Models Used
Table 7 illustrates that we are using a diverse set of models from 11 model families, combining both instruct and base versions of models. We are also covering a diverse set of sizes: from 2B up to 72B, and use both open-weights and API-based models.
Model Family Model
Yi Yi-Coder-9B
Yi-Coder-9B-Chat
Yi-Coder-1.5B-Chat
Yi-Coder-1.5B
GPT GPT-4o-mini
GPT-4o
Qwen Qwen2.5-Coder-7B
Qwen2.5-Coder-7B-Instruct
Qwen2.5-Coder-1.5B-Instruct
Qwen2.5-Coder-32B-Instruct
Qwen2.5-72B-Instruct
Qwen2.5-Coder-1.5B
Qwen2.5-Coder-14B-Instruct
Gemma codegemma-7b-it
codegemma-7b
codegemma-2b
CodeLlama CodeLlama-70b-Instruct-hf
CodeLlama-34b-Instruct-hf
CodeLlama-7b-hf
Deepseek deepseek-coder-6.7b-instruct
deepseek-coder-6.7b-base
deepseek-coder-1.3b-instruct
deepseek-coder-1.3b-base
Granite granite-8b-code-instruct-4k
granite-8b-code-base-4k
Llama Llama-3.1-8B-Instruct
Llama-3.2-3B
Llama-3.1-70B-Instruct
Llama-3.3-70B-Instruct
Llama-3.3-70B-Instruct-Turbo
Llama-3.2-1B
Llama-3.1-8B
Phi Phi-3-small-8k-instruct
Phi-3-mini-4k-instruct
phi-4
Phi-3-medium-4k-instruct
phi-2
Phi-3.5-mini-instruct
Mistral Mistral-Small-24B-Instruct-2501
StarCoder starcoder2-15B
starcoder
starcoder2-7b
starcoder2-3b
Table 7: Model families and their selected models used in DroidCollection.
Appendix B Dataset Creation and Statistics
B.1 Inverse Instructions Setup
For inverse instruction creation, we applied 4 LLMs: GPT-4o-mini, Llama3.1 8B, Qwen2.5 7B, and Phi-3 small (7B). These models were given the code, and they were asked to generate their summary and a prompt which could result in an LLM generating it. The prompt is given on LABEL:lst:code-analysis.
Listing 1: Prompt for code analysis and LLM prompt generation
# Code Analysis and LLM Prompt Generation
You are an experienced software engineer using ‘{language}‘ programming language skilled in analyzing, summarizing, and writing code. When provided with code, you break it down into its constituent parts, summarize its functionality concisely, and create prompts to guide an LLM in replicating similar outputs.
## Your Tasks:
1. **Code Summary**: Analyze the given code and summarize its purpose, logic, and functionality. Enclose this summary within [SUMMARY] and [/SUMMARY] tags.
2. **Prompt Creation**: Write a clear and specific LLM prompt that, if provided to a language model, would generate code with similar functionality and structure. Enclose the LLM prompt within [LLM_PROMPT] and [/LLM_PROMPT] tags.
Interaction will be in the following way:
### INPUT:
[CODE]
{{code}}
[/CODE]
### OUTPUT:
[SUMMARY]
{{summary}}
[/SUMMARY]
[LLM_PROMPT]
{{prompt}}
[/LLM_PROMPT]
Examples of codes, and corresponding inverse instructions are in LABEL:tab:inst_code_python, LABEL:tab:inst_code_java and LABEL:tab:inst_code_go.
B.2 DroidCollection-Personas creation
To generate DroidCollection-Personas, we started by identifying the main characteristics of a programmer. Our final list contains 9 features: Primary Programming Language, Preferred Frameworks, Field of Work, Code Commenting Style, Error-Proneness, Debugging Strategies, Code Aesthetics, Documentation Habits, Function Length Preference. The possible values for each feature are listed in Table 8.
Then we did a Cartesian product to combine all the possible combinations of these properties, and started generating the tasks, which could be performed by this programmer. For task generation, we used the GPT-4o model, and prompted it in the way shown in LABEL:lst:persona-prompt.
Listing 2: Prompt for Persona’s task generation
I have the following description of a programmer:
{description}
Write a non-trivial programming task
which matches what this person probably does at work,
you can ignore some of the persons traits. Return only the task.
After the tasks were generated, we deduplicated them using MinHash with the same parameters as for the dataset filtering. After that, the resulting tasks were used for code generation.
Property Name Values / Options
Primary Programming Language Python, Java, JavaScript, PHP, C, C#, C++, Go, Ruby, Rust
Field of Work Web Development, AI/ML, Game Development, System Programming, Embedded Systems, Data Engineering, Research, Distributed Systems Developer, IoT
Code Commenting Style Concise, Detailed, Minimal
Error-Proneness High, Medium, Low
Debugging Strategies Print Statements, Debugger, Logging
Code Aesthetics Highly Readable, Functional, Minimalist, Hard to Comprehend
Documentation Habits Detailed, Minimal, Occasional
Function Length Preference Short, Medium, Long
Table 8: List of attributes and characteristics in DroidCollection-Personas.
B.3 Dataset Statistics
In this section, we present key statistics of our dataset and compare them with existing alternatives. As shown in Table 9, our dataset includes a broader class distribution and shows greater diversity in code structure, as reflected by higher AST depth percentiles and longer line lengths. It suggests that our dataset captures more complex and varied code patterns, making it a more challenging and real-life-oriented benchmark for evaluating AI-generated code detection models. The importance of varying code lengths and difficulties is also shown in Section D.1. We also show the number of samples per generator, and programming language (not considering the datasets with <=<=< = 2 languages or generators). Several qualitative examples of samples belonging to different classes in our dataset are shown in LABEL:tab:code_samples_1 and LABEL:tab:code_samples_2.
Metric CoDet-M4 CodeGPTSensor GptSniffer DroidCollection
AST@75 15.00 12.00 15.00 15.00
AST@90 18.00 15.00 18.00 18.00
AST@99 23.00 20.00 23.15 25.00
Line@75 90.00 93.00 99.00 107.00
Line@90 113.00 112.00 117.00 135.00
Line@99 228.00 169.00 153.60 314.00
Class Distribution AI - 50% AI - 50% AI - 90% AI - 25%
Human - 50% Human - 50% Human - 10% Human - 47%
Refined - 13%
Adv. - 15%
Avg. # of samples per language 166,850 - - 148,491
Avg. # of samples per generator 50,866 - - 8,458
Table 9: Comparison of AST depth percentile, line length percentile, class distribution, and average samples per language/generator between DroidCollection and existing datasets.
Model Truncation Length
128 256 512
GptSniffer 57.05 57.20 56.64
M4 59.69 53.10 51.13
CoDet-M4 72.28 70.62 61.68
DroidDetect-Base 91.90 96.25 99.18
DroidDetect-Large 94.91 98.31 99.25
Table 10: Impact of input length truncation (measured using the ModernBERT tokeniser) on weighted F1-scores for binary classification. The most competitive numbers are highlighted in bold.
Appendix C Detailed Architectural Ablations
C.1 GCN Experiments
We used a simple 4-layer Graph Convolutional Network (GCN) to evaluate how effectively a GCN can capture structural and semantic features of code. As input, we utilised AST representations of the code, treating them as graphs. To assess the impact of node-level information, we experimented with three types of node features:
  • Dummy features – no meaningful features were provided at the node level;
  • One-hot encoded node types – encoding the syntactic type of each AST node;
  • Node content embeddings – textual embeddings derived from the string content of each node. To reduce computational overhead, we used the HashingVectorizer, which converts strings into sparse vectors by hashing tokens to fixed-dimensional indices without maintaining a vocabulary in memory.
As shown in Table 11, features based on the textual content of the node yielded the best performance, showing that the semantic information is important in distinguishing between human-written and AI-generated code.
Features 2-class 3-class 4-class
Dummy 60.02 39.27 34.17
Node Type 50.12 39.54 33.12
Text 76.67 59.10 51.14
Table 11: Comparison of different feature types used as node-level features in a GCN, based on the weighted F1-score on the validation set. The most competitive numbers are highlighted in bold.
C.2 CatBoost Experiments
Figure 1: Feature importances
Following the experimental methodology of Idialu et al. (2024) and Orel et al. (2025), we computed 733 statistical features that capture various structural properties of code. These include metrics such as the density of specific AST node types, average line length, whitespace ratio, and the number of empty lines, code maintainability index, among others.
These features were used to train CatBoost classifiers with automatically tuned hyperparameters. Figure 1 shows the top unique features ranked by SHAP (SHapley Additive exPlanations) values (Lundberg and Lee, 2017). Interestingly, the most informative features vary across the 2-, 3-, and 4-class classification tasks, suggesting that different granularities of classification are dependent on different aspects of code structure.
Nonetheless, some patterns persist across all setups. In particular, features related to the length of identifiers (variable names) and the density of comments consistently present as strong indicators for distinguishing AI-generated/Refined from human-written code.
Figure 2: Weighted F1-score comparison between models with and without fusion
C.3 Does Structure-Based Late-Fusion Improve Robustness?
To decide whether fusion is helpful for improving the detection, we combined the GCN from Section C.1 with our text-only classifier using early fusion of embeddings. We used OOD-based generalisation, and compared how well the models perform for 2, 3, and 4-class classification in OOD settings (since when trained directly, it is hard to measure the significance of the performance difference), and then compared in which scenarios each method provides a better weighted F1-score. Table 12 shows that there is no clear trend of one approach being better than another: in the binary classification task, there are more ties, fusion has a higher win-rate in 4-class classification, while the model without fusion performs best in the 3-class case. Then we compared how the difference in F1-scores between models compares to the interquartile range within the model’s predictions. As shown in Figure 2, the interquartile range is much larger than the model difference, so both models with and without fusion perform nearly equally.
Classification Tie With Without
(%) Fusion (%) Fusion (%)
2-Class 60.0 40.0 0.0
3-Class 40.0 20.0 40.0
4-Class 20.0 60.0 20.0
Table 12: Comparative task-level win-rates of DroidDetect with and without GCN late-fusion aggregated over OOD classification tasks.
Appendix D DroidDetect Stress Tests
D.1 Input Length Stress Tests
Table 10 shows that while other approaches seem to work best on short code snippets, probably because they were trained on shorter code samples (mainly functions) as shown in Table 9, our models actually get better with longer code sequences. This matters because real code is not usually just a few lines long. Another important thing is how stable our models remain across different input lengths. When we cut the input from 512 to 128 tokens, DroidDetect-Base only drops 7.28 F1-score points (from 99.18 to 91.90), and DroidDetect-Large drops just 4.34 points (from 99.25 to 94.91). This consistency suggests the generalisability of our models to various inputs.
D.2 Additional OOD Stress Testing
To evaluate the generalisation ability of our models, we tested them on additional open-source datasets containing AI-generated code. Specifically, we sampled 15,000 examples from the Swallow-Code dataset (Fujii et al., 2025), a high-quality collection of Python code from The Stack v2 (Lozhkov et al., ) synthetically refined by LLaMA3.3-70B-Instruct model. This dataset was concurrently released with our work and is highly unlikely to be part of the training distribution of any of our models, thus serving as a strong test for our models’ recall on machine-rewritten code.
We also randomly selected 15,000 samples per programming language from The Heap (Katzy et al., 2025) dataset. This dataset contains illiberally licensed code with metadata about its presence in existing code-retraining corpora. We specifically filter for samples that are not exact- or near-duplicates with any sample in major pre-training corpora (Li et al., 2023c; Lozhkov et al., ; Weber et al., 2024a). Jointly, these ensure that our curated split is extremely unlikely to be seen by models during pre-training, thus constituting a stiff test of our detectors’ recall on human-written code.
Both DroidDetect-Base and DroidDetect-Large models were tested on these datasets. On Swallow-Code, they achieved recall scores of 98.95% and 99.11% respectively. On The Heap, DroidDetect-Base reached 94.14%, while DroidDetect-Large achieved 96.28%. It shows that the models trained on our dataset can also work on other datasets robustly.
Appendix E Error Analysis and Interpretation
E.1 Error Analysis
In this section we describe and demonstrate common errors, observed in predictions of DridDetect models.
False Positives and False Negatives
Among misclassifications in the binary classification task, we observe that approximately 34% are false positives, with the remainder being false negatives. This distribution is reasonable given the larger number of negative-class samples in the dataset
Worst-Performing Language and Domain
As shown in Table 3, Research/DS domain consistently yields the lowest performance for both the Base and the Large DroidDetect models, likely due to the comparatively longer and more complex code typical in this domain. In Table 4, we observe that detecting machine-generated code in JavaScript is particularly challenging, which aligns with findings from Orel et al. (2025). We attribute this to the greater variability of coding practices found in JavaScript programs - modern frameworks of JavaScript mix coding paradigms (functional programming, OOP) - and it is also not typologically related to the rest of the languages considered in our work.
Correlation with Preserved Human-Written Code in Hybrid Cases
In hybrid scenarios, both models achieve similar F-scores for rewriting and continuation cases (91.42% and 92.47% on average). We also find a moderate negative correlation (-0.43 on average) between the proportion of preserved human-written code and the model’s F-score, indicating that the more original human code is retained, the more difficult detection becomes.
Misclassification on Adversarial Samples
As shown in Figure 5, for both the Base and Large DroidDetect models, most misclassifications occur between the AI-generated and Refined classes. Notably, adversarial samples are more frequently misclassified as human-written rather than as other machine-generated (or refined) classes. This suggests that our adversarial training strategy effectively makes these examples more human-like.
Figure 3: DroidDetect-Base classifiers confusion matrix
Figure 4: DroidDetect-Large classifiers confusion matrix
Figure 5: Confusion Matrixes for DroidDetect models
E.2 Interpretation of Predictions
We studied the patterns behind models’ predictions, using gradient attribution Ancona et al. (2018) and the attention maps of our models. Although these results are somewhat noisy and may not always be fully faithful (Laugel et al., 2019; Lyu et al., 2024), several common patterns emerge. Notably, much of the model’s attention is directed toward punctuation symbols (e.g., colons, semicolons, arrows), consistent with the findings of Clark et al. (2019), who showed that BERT-based models frequently focus on punctuation. A more detailed gradient-based analysis reveals that stylistic choices, such as long, descriptive comments and detailed explanations, typical of LLM-generated code, significantly impact the model’s decisions. Additionally, consecutive empty lines and extra spaces influence predictions, aligning with the CatBoost feature behaviour described in Section C.2. Together, these findings suggest that the structural and stylistic conventions of LLM-generated code differ from human-written code.
Examples of some misclassifications and their explanations based on gradient attribution and attention maps are given in LABEL:tab:explanations_1 and LABEL:tab:explanations_2

### v3

@@
 *[Uncaptioned image]DDroid: A Resource Suite for AI-Generated Code Detection
-Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2 Ubiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technische Universität Darmstadt {name.surname}@mbzuai.ac.ae; {name.surname}@tu-darmstadt.de
+Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2 Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE, Germany {name.surname}@mbzuai.ac.ae; {name.surname}@tu-darmstadt.de
 1 Introduction
@@
 (2024) have shown that detectors such as GPT-Zero often fail when applied to code rather than text.
-This critical observation, backed up by our experiments in Section 4, highlights the inadequacy of directly porting generic text-based models to the code domain and strongly motivates the creation of code-specific detection strategies and specialised datasets.
+This critical observation, backed up by our experiments in Section˜4, highlights the inadequacy of directly porting generic text-based models to the code domain and strongly motivates the creation of code-specific detection strategies and specialised datasets.
 Our work responds to this need by providing a large-scale, multifaceted suite specifically curated for AI-generated code, designed to foster the development and rigorous testing of detection techniques attuned to the unique characteristics of programming languages and AI-generated software.
@@
 Table 1: Comparison of DroidCollection to other AI-generated code detection datasets shows broader domain and model coverage, as well as unique inclusion of varied sampling strategies and adversarially humanized data.
-We compare our dataset with existing ones in Table 1, which shows that our dataset is not only among the largest to date, but also captures a broader range of variations.
-Additional details about key characteristics are provided in Section B.3.
+We compare our dataset with existing ones in Table˜1, which shows that our dataset is not only among the largest to date, but also captures a broader range of variations.
+Additional details about key characteristics are provided in Section˜B.3.
 In this section, we detail the curation of the human-generated, machine-generated, and machine-refined splits of DroidCollection.
-The adversarially humanised data collection is deferred to Section 3.4.
+The adversarially humanised data collection is deferred to Section˜3.4.
 3.1 Human-Authored Code Acquisition
@@
 Then, we generated code, using base and instruction-tuned LMs from 11 model release families, namely: Llama (Grattafiori et al., 2024), CodeLlama, GPT-4o, Qwen (Qwen et al., 2025), IBM Granite (Mishra et al., 2024), Yi (AI et al., 2025), DeepSeek (Guo et al., 2024), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), Mistral (AI, 2025), Starcoder (Li et al., 2023c).
-The list of generators per model family is detailed in Appendix A.
+The list of generators per model family is detailed in Appendix˜A.
 Our dataset covers three domains: general use code, algorithmic problems, and research/data-science code.
@@
 However, in cases where a summary could be extracted from the response but the instruction could not, we used the summary to re-generate the instruction.
-This experiment with details about the prompts and the models we used is illustrated in Section B.1.
+This experiment with details about the prompts and the models we used is illustrated in Section˜B.1.
 This type of generation allows us to cover a wide range of prompts, simulating a diversity of user-LM interactions, which is common in the real world.
@@
 These tasks, along with their corresponding descriptions, termed DroidCollection-Personas, were then used to generate code samples.
-More details about the DroidCollection-Personas are outlined in the Section B.2.
+More details about the DroidCollection-Personas are outlined in the Section˜B.2.
 3.3 Machine-Refined Data
@@
 Once trained, the models are used to generate new “machine-humanised” code samples.
-We filter their outputs as in Section 3.6 to keep only high-quality adversarial examples.
+We filter their outputs as in Section˜3.6 to keep only high-quality adversarial examples.
 Finally, we obtained a nearly 1:1 ratio of prompt-based vs.
@@
 (2020) that after greedy decoding, it was easier to detect AI-generated texts compared to when using other decoding techniques.
-To capture this challenging behaviour and reflect the diversity of generative systems, we experimented with various decoding strategies, as shown in Table 2.
+To capture this challenging behaviour and reflect the diversity of generative systems, we experimented with various decoding strategies, as shown in Table˜2.
 3.6 Data Filtering
@@
 It means, that there is a possibility that among human written codes there could be samples, created with the help of LLMs.
-This potential issue is tackled in Section 5.
+This potential issue is tackled in Section˜5.
 Model 2-Class 3-Class
@@
 Additionally, we train other models directly on our dataset using a multi-class objective: (i) a simple GCN(Kipf and Welling, 2017); (ii) a CatBoost classifier(Prokhorenkova et al., 2018), following procedures similar to the Whodunit paper; and (iii) two encoder-only transformers, ModernBERT-Base and ModernBERT-Large (Warner et al., 2024), denoted as DroidDetectCLS-Base and DroidDetectCLS-Large.
-Full details are provided in Sections C.1, C.2 and 5.
+Full details are provided in Sections˜C.1, C.2 and 5.
 4.2 RQ1: What is the value of extensive data collection for training robust detectors?
-Table 1 highlights a key limitation of existing datasets: they often lack diversity.
-Tables 3 and 4 show that this limitation significantly impacts detector performance in realistic settings.
+Table˜1 highlights a key limitation of existing datasets: they often lack diversity.
+Tables˜3 and 4 show that this limitation significantly impacts detector performance in realistic settings.
 Baseline detectors - the ones illustrated in Zero-Shot Baselines section of the tables - underperform on our test split compared to even simple fine-tuned baselines like GCN and CatBoost.
@@
 We selected the base version of the backbone for this restricted training scenario since it was comparable to most of the chosen baselines in terms of size.
-Table 4 shows that, under restricted training conditions, models tend to generalise better to syntactically similar languages.
+Table˜4 shows that, under restricted training conditions, models tend to generalise better to syntactically similar languages.
 For example, a model trained on C/C++ performs reasonably well on C# and Java.
 However, for typologically isolated languages such as Python or JavaScript, all models not trained specifically for it tend to struggle in this setting.
-Table 3 illustrates that the models trained on a single domain have a high discrepancy in scores for other domains.
+Table˜3 illustrates that the models trained on a single domain have a high discrepancy in scores for other domains.
 For example, as can be seen in the OOD Evaluation section, the classifiers trained on algorithmic problems suffer with the general codes, and show comparable low performance on Research/DS codes, and vice versa.
 4.4 RQ3: How robust are models to adversarial samples?
-Finally, we test adversarial robustness using challenging samples designed to evade detection (Table 5).
+Finally, we test adversarial robustness using challenging samples designed to evade detection (Table˜5).
 Here, many baseline detectors struggle: for instance, GPT-Zero achieves only 0.10 recall.
@@
 FastDetectGPT GPTSniffer M4 CoDet-M4 GPT-Zero M4FT GPT-SnifferFT CoDet-M4FT DroidDetectCLS-Base DroidDetectCLS-Large
-Human-written 0.84 0.65 0.40 0.38 0.53 0.91 0.97 0.96 0.93 0.98
+[…]
 Adversarial samples 0.48 0.49 0.73 0.63 0.10 0.67 0.55 0.51 0.92 0.92
@@
 Specifically, we trained a 4-layer Graph Convolutional Network over the AST representation of codes to evaluate its ability to distinguish AI-generated from human-written code.
-The results are shown in Section C.1.
+The results are shown in Section˜C.1.
 We can see that while structural signals are informative, GCNs alone are not sufficient to achieve strong generalisation.
@@
 This model was selected for the inference efficiency (Warner et al., 2024), and suitability for code-related tasks.
-However, as shown in Section C.3, this fusion strategy yielded only a marginal improvement.
+However, as shown in Section˜C.3, this fusion strategy yielded only a marginal improvement.
 Consequently, we decided to use a text-only encoder for the final model.
@@
 We refer to these models as DroidDetectSCL.
-Table 6 demonstrates the small but consistent performance gain unlocked using metric learning.
+Table˜6 demonstrates the small but consistent performance gain unlocked using metric learning.
 Finally, we addressed the problem of noisy and mislabelled training data.
@@
 This manner of self-bootstrapping datasets has an extensive track record in image (Yalniz et al., 2019; Xie et al., 2020) and text (Wang et al., 2022) representation learning, relying on the tendency of neural networks to understand patterns in clean labels before they overfit to noisy data (Feldman and Zhang, 2020).
-Incorporating this filtering into our training yielded our final DroidDetect models, which, as shown in Table 6, perform the best across the classification settings.
+Incorporating this filtering into our training yielded our final DroidDetect models, which, as shown in Table˜6, perform the best across the classification settings.
 We trained all models for 3 epochs, using AdamW (Loshchilov and Hutter, 2019) optimiser, setting the top learning rate to 5e-5, and applying the linear warmup (proportion 0.1) with cosine decay learning rate scheduler.
@@
 We advocate for the responsible use in strictly legitimate research and educational contexts.
+Acknowledgements
+At TU Darmstadt, this research work has been supported by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE and by HUAWEI Technologies (Ireland) Co., Ltd.
 Appendix A List Of Models Used
-Table 7 illustrates that we are using a diverse set of models from 11 model families, combining both instruct and base versions of models.
+Table˜7 illustrates that we are using a diverse set of models from 11 model families, combining both instruct and base versions of models.
 We are also covering a diverse set of sizes: from 2B up to 72B, and use both open-weights and API-based models.
@@
 Our final list contains 9 features: Primary Programming Language, Preferred Frameworks, Field of Work, Code Commenting Style, Error-Proneness, Debugging Strategies, Code Aesthetics, Documentation Habits, Function Length Preference.
-The possible values for each feature are listed in Table 8.
+The possible values for each feature are listed in Table˜8.
 Then we did a Cartesian product to combine all the possible combinations of these properties, and started generating the tasks, which could be performed by this programmer.
@@
 In this section, we present key statistics of our dataset and compare them with existing alternatives.
-As shown in Table 9, our dataset includes a broader class distribution and shows greater diversity in code structure, as reflected by higher AST depth percentiles and longer line lengths.
+As shown in Table˜9, our dataset includes a broader class distribution and shows greater diversity in code structure, as reflected by higher AST depth percentiles and longer line lengths.
 It suggests that our dataset captures more complex and varied code patterns, making it a more challenging and real-life-oriented benchmark for evaluating AI-generated code detection models.
-The importance of varying code lengths and difficulties is also shown in Section D.1.
-We also show the number of samples per generator, and programming language (not considering the datasets with <=<=< = 2 languages or generators).
+The importance of varying code lengths and difficulties is also shown in Section˜D.1.
+We also show the number of samples per generator, and programming language (not considering the datasets with <=<= 2 languages or generators).
 Several qualitative examples of samples belonging to different classes in our dataset are shown in LABEL:tab:code_samples_1 and LABEL:tab:code_samples_2.
@@
 To reduce computational overhead, we used the HashingVectorizer, which converts strings into sparse vectors by hashing tokens to fixed-dimensional indices without maintaining a vocabulary in memory.
-As shown in Table 11, features based on the textual content of the node yielded the best performance, showing that the semantic information is important in distinguishing between human-written and AI-generated code.
+As shown in Table˜11, features based on the textual content of the node yielded the best performance, showing that the semantic information is important in distinguishing between human-written and AI-generated code.
 Features 2-class 3-class 4-class
@@
 These features were used to train CatBoost classifiers with automatically tuned hyperparameters.
-Figure 1 shows the top unique features ranked by SHAP (SHapley Additive exPlanations) values (Lundberg and Lee, 2017).
+Figure˜1 shows the top unique features ranked by SHAP (SHapley Additive exPlanations) values (Lundberg and Lee, 2017).
 Interestingly, the most informative features vary across the 2-, 3-, and 4-class classification tasks, suggesting that different granularities of classification are dependent on different aspects of code structure.
@@
 C.3 Does Structure-Based Late-Fusion Improve Robustness?
-To decide whether fusion is helpful for improving the detection, we combined the GCN from Section C.1 with our text-only classifier using early fusion of embeddings.
+To decide whether fusion is helpful for improving the detection, we combined the GCN from Section˜C.1 with our text-only classifier using early fusion of embeddings.
 We used OOD-based generalisation, and compared how well the models perform for 2, 3, and 4-class classification in OOD settings (since when trained directly, it is hard to measure the significance of the performance difference), and then compared in which scenarios each method provides a better weighted F1-score.
-Table 12 shows that there is no clear trend of one approach being better than another: in the binary classification task, there are more ties, fusion has a higher win-rate in 4-class classification, while the model without fusion performs best in the 3-class case.
+Table˜12 shows that there is no clear trend of one approach being better than another: in the binary classification task, there are more ties, fusion has a higher win-rate in 4-class classification, while the model without fusion performs best in the 3-class case.
 Then we compared how the difference in F1-scores between models compares to the interquartile range within the model’s predictions.
-As shown in Figure 2, the interquartile range is much larger than the model difference, so both models with and without fusion perform nearly equally.
+As shown in Figure˜2, the interquartile range is much larger than the model difference, so both models with and without fusion perform nearly equally.
 Classification Tie With Without
@@
 D.1 Input Length Stress Tests
-Table 10 shows that while other approaches seem to work best on short code snippets, probably because they were trained on shorter code samples (mainly functions) as shown in Table 9, our models actually get better with longer code sequences.
+Table˜10 shows that while other approaches seem to work best on short code snippets, probably because they were trained on shorter code samples (mainly functions) as shown in Table˜9, our models actually get better with longer code sequences.
 This matters because real code is not usually just a few lines long.
@@
 Worst-Performing Language and Domain
-As shown in Table 3, Research/DS domain consistently yields the lowest performance for both the Base and the Large DroidDetect models, likely due to the comparatively longer and more complex code typical in this domain.
-In Table 4, we observe that detecting machine-generated code in JavaScript is particularly challenging, which aligns with findings from Orel et al.
+As shown in Table˜3, Research/DS domain consistently yields the lowest performance for both the Base and the Large DroidDetect models, likely due to the comparatively longer and more complex code typical in this domain.
+In Table˜4, we observe that detecting machine-generated code in JavaScript is particularly challenging, which aligns with findings from Orel et al.
 (2025).
@@
 Misclassification on Adversarial Samples
-As shown in Figure 5, for both the Base and Large DroidDetect models, most misclassifications occur between the AI-generated and Refined classes.
+As shown in Figure˜5, for both the Base and Large DroidDetect models, most misclassifications occur between the AI-generated and Refined classes.
 Notably, adversarial samples are more frequently misclassified as human-written rather than as other machine-generated (or refined) classes.
@@
 A more detailed gradient-based analysis reveals that stylistic choices, such as long, descriptive comments and detailed explanations, typical of LLM-generated code, significantly impact the model’s decisions.
-Additionally, consecutive empty lines and extra spaces influence predictions, aligning with the CatBoost feature behaviour described in Section C.2.
+Additionally, consecutive empty lines and extra spaces influence predictions, aligning with the CatBoost feature behaviour described in Section˜C.2.
 Together, these findings suggest that the structural and stylistic conventions of LLM-generated code differ from human-written code.